OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 2) and a directrix at y = -2.

OpenStudy (anonymous):

I got: (x-2)^2 = 8(y-0) which simplifies to x^2-4x+4=8y but that is not any of my options...

OpenStudy (anonymous):

My options: y^2 = 2x y = 1/2*x2 y^2 = 8x y = 1/8*x2

OpenStudy (wolf1728):

I'd say the answer is y = 1/8*x2

OpenStudy (triciaal):